### v3

# Doing Experiments and Revising Rules with Natural Language and Probabilistic Reasoning

Abstract
We build a computational model of how humans actively infer hidden rules by doing experiments. The basic principles behind the model is that, even if the rule is deterministic, the learner considers a broader space of fuzzy probabilistic rules, which it represents in natural language, and updates its hypotheses online after each experiment according to approximately Bayesian principles. In the same framework we also model experiment design according to information-theoretic criteria. We find that the combination of these three principles—explicit hypotheses, probabilistic rules, and online updates—can explain human performance on a Zendo-style task, and that removing any of these components leaves the model unable to account for the data.
Keywords: induction; LLM; active learning
Introduction
An important way that humans grow their knowledge of the world is by experimentation, intervention, and other forms of active learning. This process is most clearly present in the experimental sciences, but similar processes of active inference contribute to human learning starting even in infancy \@BBOPcitep\@BAP\@BBN(Mendez et al., 2023; Cook et al., 2011; Carey, 1985; Schulz, 2012; Gopnik et al., 1999)\@BBCP. Within everyday adult learning and thinking, active experimentation helps us quickly learn to use new devices, web pages, and other tools.
In this work, we aim to give a computational account of how people propose experiments and revise their beliefs. We work within the context of a game called Zendo, which is meant to be a microcosm of experimental science. In Zendo, there is a secret logical rule which determines whether an assembly of colored blocks is a positive example of a hidden concept (referred to as ‘having the zendo nature’). Players try to discover this rule by building structures of colored blocks and asking if their structure obeys the hidden logical rule. Each structure they build is analogous to an experiment designed to probe the rule that determines if something ‘has the Zendo nature’. Players both try to infer the rule and predict which new structures conform to the rule. Studying the experiments and judgments that humans produce across a variety of different hidden rules gives a rich domain for modeling, and we will use human data from a study of Zendo \@BBOPcitep\@BAP\@BBN(N. Bramley et al., 2018)\@BBCP in which to test and develop different models of how humans propose experiments and update their beliefs about hidden rules and regularities (Fig. 1).
Figure 1: Illustration from \@BBOPcite\@BAP\@BBNN. Bramley et al. (2018)\@BBCP of their Zendo setup, which we use wholesale. Human participants build structures out of colored cones of different sizes (left panel). Each structure is an experiment designed to help the participant learn which structures have the ‘Zendo nature,’ which this experiment equivalently described as ‘Emitting zeta waves.’ After doing 8 experiments and receiving feedback on which structures follow the hidden rule (top right panel), participants predicted which new structures also obeyed the rule (bottom right panel).
Our model will start with simple core principles: Learners represent hypotheses in an inner symbolic language-of-thought \@BBOPcitep\@BAP\@BBN(Fodor, 1975; Piantadosi et al., 2016)\@BBCP, update their beliefs in approximately Bayesian ways \@BBOPcitep\@BAP\@BBN(Chater & Oaksford, 2008)\@BBCP, and (try to) propose experiments that are optimal in an information-theoretic sense \@BBOPcitep\@BAP\@BBN(Settles, 2009)\@BBCP. Informally, having Bayesian beliefs means that the learner attaches a probability to every possible rule, in light of all of the experiments that it has seen. Optimal experiments are those which are most likely to cause large changes to these beliefs. Within the context of zendo, this means an ideal Bayesian experimenter would keep track of every possible rule (of which there are infinitely many), consider every single experiment (again, infinite), and then compute the expected belief change for each experiment to decide what to try next (each of which involves considering infinitely many hypotheses).
Actual learners are bounded, and can only consider a small number of hypotheses at a time. How then can they effectively approximate the above ideal probabilistic reasoner? Our model uses two tricks, both of which we believe have analogues in human thinking. First, our model considers fuzzy probabilistic rules, despite knowing that the true rule is deterministic. A probabilistic rule is equivalent to a distribution over deterministic ones, so a small collection of fuzzy rules can stand in for a broader space of deterministic hypotheses. Second, we assume that the learner generates a small number of candidate hypotheses using a heuristic data-driven bottom-up mechanism \@BBOPcitep\@BAP\@BBN(Ullman et al., 2016)\@BBCP, which are then re-analyzed in a more properly Bayesian manner. This bottom-up proposal process helps focus the learner on the subset of the hypothesis space which is actually likely, instead of wasting computational effort on low-probability regions of the hypothesis space.
Our model instantiates several other claims about how human learners make experiments and revise beliefs. For example, the model updates its beliefs on-the-fly by revising its previous hypothesis, instead of constructing its beliefs all at once in a batch-style update, known as online inference. Prior work \@BBOPcitep\@BAP\@BBN(Thaker et al., 2017)\@BBCP has found that online inference helps explain ordering effects such as anchoring \@BBOPcitep\@BAP\@BBN(Kahneman, 2011; MacDonald, 1994)\@BBCP. We also model the language-of-thought (the encoding of hypotheses) as natural-language English text, which allows pretrained large language models (LLMs) to act as bottom-up data-driven hypothesis generators. Recent work has found that LLM-generated natural language hypotheses can predict certain judgments of human learners, particularly when embedded in a Bayesian framework \@BBOPcitep\@BAP\@BBN(Ellis, 2023)\@BBCP.
Our model accounts for the whole life cycle of a Zendo game, starting from initial hypothesis formation, sequential experimentation, and predicting which members of a test set ‘have the zendo nature’. In other words, it can actually play Zendo and therefore can be compared head-to-head against people. We find that the basic principles of the model—fuzzy rules, online inference, and bottom-up natural language hypothesis generation—are all required to accurately predict human behavioral data. This suggests that when humans function as ‘intuitive experimentalists’, they might do so by constructing a small number of fuzzy hypotheses, and then reasoning about them in ways that roughly accord with the norms of probabilistic inference.
Zendo Domain and Human Data
Zendo is a game where a player seeks to infer a hidden binary rule about assemblies of colored blocks. The game starts by providing the player with a positive scene that follows the hidden rule. Then, the player queries an oracle as to a particular scene follows the rule or not, or makes a guess about the secret rule. The game ends when they guess correctly.
\@BBOPcitet\@BAP\@BBNN. Bramley et al. (2018)\@BBCP introduces a 2D version of the Zendo game shown in Figure 1. The scenes consist of blocks, each with its own color (red, blue, green) and size (small, medium, large). The blocks can have different orientations and positions in a 2D scene. They may and may not touch each other. The game starts with an initial phase where a rule-following scene is given, followed by 7 rounds of active learning phase where the player gets to query an oracle for ground truth prediction. At the end, the player enters a prediction phase where they are asked to give predictions for 8 test scenes (4 rule-following and 4 not rule-following). \@BBOPcitet\@BAP\@BBNN. Bramley et al. (2018)\@BBCP study human gameplay on 10 rules, collecting data from 30 participants who each play Zendo 10 times (once per rule). They use a cover story of an alien planet where some arrangements of blocks emit radiation, and the task is to figure out a rule predicting radiation emission.
Zendo is most naturally framed as a visual-physical concept learning problem. For our model, however, we will work with discrete symbolic descriptions of scenes. This makes that problem more compatible with the language-of-thought paradigm, and also allows using LLMs to operationalize the language of thought. We therefore modify \@BBOPcitet\@BAP\@BBNN. Bramley et al. (2018)\@BBCP’s version of Zendo by associating each block in a scene with discrete attributes instead. The five attributes are color (red, blue, green), size (small, medium, large), orientation (upright, left, right, strange), groundedness (grounded, ungrounded), and touching (which blocks it touches). While this natural language version of the game removes continuous attributes, such as $x,y$ position and orientation in 2D space, from its scene representations, these five attributes still maintain the complexity of the game and are sufficient for playing Zendo with 9 out of their 10 rules. [^1] The 9 rules we use are “there’s a red”, “all are the same size”, “nothing is upright”, “one is blue”, “there’s a small blue”, “all are blue or small”, “a red is bigger than all non reds”, “some touch”, “a blue and a red touch”, and “some pieces are stacked”. The excluded rule is “some pieces are stacked”, as it requires reasoning about the $x$-position of blocks.
Several prior works  \@BBOPcitep\@BAP\@BBN(N. Bramley et al., 2018; Fränken et al., 2022; N. R. Bramley & Xu, 2023)\@BBCP have provided computational models that explain aspects of how humans play Zendo. Within this literature, we contribute both new modeling approaches and also a system that can play Zendo end-to-end. We also provide the finding that online inference and fuzzy rules can offer good accounts of human Zendo play, and provide a working demonstration of how pretrained knowledge from language models can guide efficient inference and active experimentation.
Model
Probabilistic Model
In order to have a unified mathematical language to describe both forming hypotheses and doing experiments, we define a model that includes natural-language hypotheses $h\in\Sigma^{*}$, experiments (Zendo scenes) $x\in\mathcal{X}$, and feedback $y\in\{0,1\}$ (whether $x$ follows the rule). We also equip $h$ with a real-valued parameter $\theta$ which controls the ‘fuzziness’ of the rule $h$. As new experiments are proposed sequentially, we index experiments and feedback with subscripts, i.e. $x_{t}$ and $y_{t}$ for the $t^{\text{th}}$ experiment and feedback, respectively.
We define a joint probability distribution over hypotheses and feedback, given experiments, as
$p(h,y_{1:T},\theta|x_{1:T})=p(h)p(\theta)\prod_{1\leq t\leq T}p(y_{t}|x_{t},h,\theta)$
In other words, we assume that there is a latent rule $h$ that existed prior to any experiments, with a fixed fuzziness $\theta$, and that each label $y_{t}$ is a result of rule $h$ and the experiment $x_{t}$. The prior beliefs about $h$, written $p(h)$, are defined to be inversely proportional to the number of words in $h$, which gives a bias toward shorter (simpler) hypotheses:
$p(h)\propto\frac{1}{word\_count(h)}$
The choice of $p(y|x,h)$ is one of the main ingredients of our model. Recent work on inductive reasoning with LLM \@BBOPcitep\@BAP\@BBN(Wang et al., 2023; Qiu et al., 2023)\@BBCP opt for a deterministic likelihood model, meaning $p(y|x,h)\in\{0,1\}$. Though a deterministic likelihood may seems natural, given that $h$ is a “rule”, we observe, and will show that a hard all-or-none likelihood model does not give a good fit to human behavior within Zendo.
We choose the likelihood based on the idea that while the secret rule of a Zendo game is always deterministic, humans might internally represent probabilistic rules, and might also assume that some of the examples are occasionally mislabeled. Then, we define $p(y|x,h,\theta)$ as follows:
$p(y=1|x,h,\theta)=\;\left[\begin{array}[]{cc}\theta\text{, if }f_{h}(x)=1\\ \epsilon\text{, otherwise}\end{array}\right]$
where $\epsilon$ is a noise parameter modeling the small probability of an example being mislabeled, and $f_{h}:\mathcal{X}\rightarrow\{0,1\}$ means that $x$ obeys the rule $h$. We set $\epsilon=0.1$, and we evaluate $f_{h}(x)$ by prompting an LLM to translate $h$ into Python code, which can then be run on $x$.
[…]
Having now defined the mathematical components of the model, at least in the abstract, we can now define the computational problems that the learner must solve. First, the learner must form probabilistic posterior beliefs about the rule (and its fuzziness), following Bayes:
$p(h,\theta|x_{1:T},y_{1:T})$ $\propto p(h)p(\theta)\prod_{1\leq t\leq T}p(y_{t}|x_{t},h,\theta)$
Given a test structure $x_{\text{test}}$, an ideal learner uses the beliefs above to predict a label $y_{\text{test}}$ following:
$p(y_{\text{test}}$ $|x_{\text{test}},x_{1:T},y_{1:T})=$
$\sum_{h}\int p(y_{\text{test}}|h,\theta,x_{\text{test}})p(h,\theta|x_{1:T},y_{1:T})\text{d}\theta$
For the active learning phase, an ideal learner conducts an optimal experimentation, which we will give more intuition for later, defined as:
$x^{*}=\operatorname*{arg\,max}_{x\in\mathcal{X}}\mathop{\mathbb{E}}_{p(y|x_{1:T},y_{1:T},x)}[D_{\mathrm{KL}}(p(h|x_{1:T},y_{1:T},x,y)||p(h|x_{1:T},y_{1:T}))]$
The above computations are intractable because they involve considering the infinitely large set of all hypotheses and experiments. We will describe approximation methods in the following sections.
Revising Rules: Online Inference
Exact inference is not computationally tractable when the hypothesis space is natural language. We introduce here a Sequential Monte Carlo (SMC) algorithm \@BBOPcitep\@BAP\@BBN(Doucet et al., 2001)\@BBCP, which tracks a small pool of hypotheses that evolve over time as new experiments are performed. This helps keep track of those hypotheses that have high posterior, allowing us to perform approximate predictions tractably. This part of our model should be seen within the tradition of using SMC methods to give mechanistic, process-level accounts of online learning \@BBOPcitep\@BAP\@BBN(Griffiths et al., 2015a)\@BBCP. Fig. 2 gives an overview of how this process works, and the caption gives an in-depth walkthrough of a single iteration of the algorithm: after each experiment, hypotheses are reweighed in light of the new data, revised to better explain the latest experiment, and last, a resampling stage multiplies high probability hypotheses while eliminating low-likelihood hypotheses.
Figure 2: Sequential Monte Carlo method tracks a small number of hypotheses (called particles), each of which is a natural language rule. Upon performing a new experiment, the particles are reweighed according to how well they can explain this new data, and then revised using a LLM to better fit the outcome of the latest experiment (“rejuvenated”). There then stochastically resampled—which kills off low-probability hypotheses—and the process repeats for another iteration. This figure shows one full iteration of SMC when a scene with a large red touching a medium blue is observed to be positive (has the zendo nature).
At the beginning of the algorithm, when the first data point is observed, we sample hypotheses $h_{1}^{(1)},...,h_{1}^{(n)}$ from an initial proposal distribution $q(h)$ defined by an LLM. The hypotheses are treated as particles, and $n$ is the number of particles. For simplicity, we assume $q(h|x,y)$ is uniform, so we assign particle weights $w_{1}^{(1)},...,w_{1}^{(n)}$ to be just the prior probability of each hypothesis. The $\theta_{1}^{(1)},...,\theta_{1}^{(n)}$ associated with each particle were set to their MAP estimate. A resampling technique is applied to normalize the weights at the end.
Then, as a new observation $x_{t},y_{t}$ comes in from the active learning phase, our algorithm first performs reweighting, then a rejuvenation step (revising rules), and lastly resampling. Reweighting involves multiplying $w^{(i)}$ with $p(y_{t}|x_{t},h_{t}^{(i)},\theta_{t}^{(i)})$ to incorporate the new observation and additionally with $\frac{p(X,Y,h_{t}^{(i)},\theta_{t}^{(i)})}{p(X,Y,h_{t}^{(i)},\theta_{t-1}^{(i)})}$ to account for the changing from $\theta_{t-1}^{(i)}$ to $\theta_{t}^{(i)}$.
After the reweighting step, we perform a rejuvenation step \@BBOPcitep\@BAP\@BBN(Chopin, 2002; Thaker et al., 2017)\@BBCP where we consider $k$ local moves proposed by a local proposer $q(h_{evolved}|h_{original},x_{t},y_{t})$ defined by prompting an LLM. We sample a batch of $k=15$ local moves for each of the 5 unique particles with lowest weights. Then, all particles that share the same hypothesis with these 3 unique particles are rejuvenated. Each particle only performs a move if it will get a higher joint probability after the move. If there are multiple moves that will result in higher joint probability, sample one with probability proportional to the joint probabilities. We update the weight with $\frac{p(y_{t}|x_{t},h_{evolved},\theta_{evolved})}{p(y_{t}|x_{t},h_{original},\theta_{original})}$. At the end, we perform the resampling step before the start of the next iteration.
It should be noted that the effectiveness of this algorithm depends on the quality of the initial and local proposers, $q(h|x,y)$ and $q(h_{evolved}|h_{original},x,y)$. The recent success of LLMs helps enable our algorithm: a properly prompted LLM can be a phenomenal hypothesis proposer \@BBOPcitep\@BAP\@BBN(Qiu et al., 2023)\@BBCP. This is despite the well-known unreliability of LLMs, because proposal distributions do not need to be fully reliable. Instead, proposals merely need to (sometimes) generate good samples. Reweighing those samples under the probabilistic model helps correct mistakes made by the LLM, and it is fundamentally the cooperation of Bayes and the LLM that allows models like ours to track beliefs with few particles \@BBOPcitep\@BAP\@BBN(Ellis, 2023)\@BBCP. For example, in our experiments, we use only 25 particles, which we think is about the same order of magnitude as what humans might track in their head. The prompts used in our experiments can be founded in the appendix. Another observation is the algorithm, in its current form, is heuristic and does not provably target the true posterior. Modifications of the local proposer and the rejuvenation step to ensure convergence guarantees are left for future work.
Doing Experiments: Active Learning
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: Comparing human and model accuracy on test scenes after 7 rounds of experimentation. Each point is a different rule assessed either on structures that follow or do not follow the rule ($\pm$±RF). Removing even a single component of the model significantly degrades its ability to predict human accuracy.
To perform active learning, we need an experiment proposer and a measure of informativeness of an experiment. The most informative experiment, among many proposed candidates, is chosen to be conducted. In the context of Zendo, doing an experiment corresponds to building a scene out of Zendo blocks, and the outcome of the experiment is 0 or 1 depending on if the assembly has the zendo nature.
Like our probabilistic inference strategy, we will propose experiments using an LLM, bur then reassess those proposals under probabilistic criteria. Particularly, we provide an LLM with the hypotheses we are keeping track of at each iteration and ask it to generate experiments that support and falsify each of the hypotheses. Empirically, this process yields a diverse pool of experiments.
For the measure of informativeness, we opt for the standard metric commonly used in the active learning literature \@BBOPcitep\@BAP\@BBN(Settles, 2009)\@BBCP: expected model change. It is defined as follows:
$\mathop{\mathbb{E}}_{p(y|x_{1:T},y_{1:T},x)}[D_{\mathrm{KL}}(p(h|x_{1:T},y_{1:T},x,y)||p(h|x_{1:T},y_{1:T}))]$
This captures how much, in expectation, the posterior belief would change if $x$ is queried for ground truth prediction. It can be argued that an experiment is informative if it significantly changes belief, so in this sense, the expected model change measures information gain.
We use GPT-4 Turbo for hypothesis generation/revision, while GPT-3.5 translates language to code \@BBOPcitep\@BAP\@BBN(OpenAI, 2023)\@BBCP.
Results
Figure 4: Comparing human and model accuracy on the 8 test scenes after 7 rounds of experimentation with breakdowns for the 4 rule-following (RF) and 4 not rule-following (Not RF) test scenes. The rows and columns of the grid of plots are the methods and the 9 rules respectively. Online inference with fuzzy rules, presented in the last row, most closely matches the accuracy breakdown observed in humans.
[…]
For a more fine-grained analysis, we noticed that across many rules, there was a significant difference in human accuracy on RF and Not RF test structures, with which condition people found easier depending on the underlying rule. One possible explanation is that, for many rules, it is easy for human to learn what doesn’t follow the rule, but very hard to learn what does follow the rule (or vice versa). Fig. 4 illustrates this and compares it against what each model thinks should be the easier condition. We again see that the full model (online learning of fuzzy rules) can successfully predict the direction of most of these trends, unlike its lesioned versions. This suggests the model might be able to capture trends both across rules and also within rules.
Discussion
Active Inference as a Cognitive Testbed
We found that our model was surprisingly fragile: Removing a single piece causes it to not merely degrade in fit, but to fail utterly. We suspect that this may be because of the feedback loop between inference and experimentation that occurs during Zendo—and indeed, which occurs during human development, and in the history of science. Subpar inference causes poor experiments which further degrades inference, creating a cascade of failures. Therefore, studying the interplay of experimentation and hypothesis creation may be an especially sensitive testbed in which to study the basics of how people build new knowledge by experimenting with the world.
Bounded Rationality and Approximate Inference
While we define the computational quantities a learner wants to calculate in the previous section, it would be misleading to think that achieving exact inference of these quantifies would result in a more human-like model. The theory of bounded rationality \@BBOPcitep\@BAP\@BBN(Simon, 1955)\@BBCP argue that humans do not always perform the most rational action due to their limited cognitive resources. Building on this concept of bounded rationality, research on resource-limited analysis \@BBOPcitep\@BAP\@BBN(Griffiths et al., 2012, 2015b; Lieder & Griffiths, 2020)\@BBCP suggest that approximate inference algorithms may provide a better account for human cognition. Our use of approximate inference model can then be viewed as an implementation of these frameworks.
Why probabilistic beliefs are necessary
One might ask if probabilistic beliefs are necessary: can we simply track of a small list of hypotheses without attaching probabilities to them, or even just track the single most likely hypothesis? We need probabilistic beliefs because they are fundamental to experiment design: Good experiments tease apart competing hypotheses, and the best contenders are those with high posterior probability. Just tracking a single hypothesis would not give any guidance on what experiment to do next, and failing to track the high-probability regions of the posterior would end up performing experiments that test many bad hypotheses while leaving the more likely ones uncontested. Consequently, the learner would be unable to narrow down the correct rule.
Limitations and open directions
The work presented is limited in important ways that suggest the next steps. Most immediately the hypotheses considered by our work are simple and stereotyped in form. Much of the promise of using natural language, or another language of thought, is that it exposes a rich, expressive representation for combining and creating new ideas. Hypothesis generation, both in science and in everyday thinking, often involves conjecturing the existence of unseen objects, not just unknown regularities. Consider for instance Mendel’s discovery of alleles, or conjecturing the existence of an unknown intruder upon discovering footprints in one’s house. These forms of abductive thinking, which often go hand-in-hand with hypothesis formation, are absent from our model.
The model as an experimental scientist
The work here is most directly an account of human behavior within the context of the game Zendo. However, our model is also more broadly inspired by the mental activities of experimental scientists as they build theories and models, weigh hypotheses, and design experiments. Two basic features of our approach reflect scientific experimentation and theory building. First, scientific theories apply only within a particular regime. For example, many equations in physics only apply when objects are moving slowly, and many thermodynamic equations only apply at equilibrium. Outside this regime, the theory no longer applies, and no longer makes clear predictions. Similarly, our model hypothesizes fuzzy probabilistic rules which can either predict a category with high confidence, or outside the regime in which the rule applies, they can fail to make a decisive prediction.
The second way in which this work reflects the practices of experimental science is that it builds its hypotheses via an incremental evolutionary process. In this way, our model is best thought of as performing what is sometimes called ‘normal science’ — where one works within an existing paradigm, and considers piecemeal evidence — and does not model paradigm shifts \@BBOPcitep\@BAP\@BBN(Kuhn, 1962)\@BBCP (what a scientific theorist might pursue) or deeper conceptual changes \@BBOPcitep\@BAP\@BBN(Carey, 1985)\@BBCP (what happens during child development), both of which require deep reanalysis of ua broad batch of past data.
Appendix A Appendix
Prompts
The prompts used in all of our experiments can be found in Tables 1, 2, 3 and 4. In Table 1, we engineer the prompts for online inference so that they only output simple rules; this approach helps the proposer output hypotheses with higher priors, since our prior is defined by the number of words in the rule. We cannot apply this trick to batch inference because, unlike online inference, it does not evolve simpler rules into more complex ones.
Batch Inference (with and without Rule Refinement) with Hard Rules {spverbatim} Given the following structures described with att_summary of blocks in the structures: text_c Please list num possible rules about the attributes in a structure that differentiate the good structures from the bad structures. Keep in mind that 1. All bad structures must violate the rules. 2. Orders of blocks in a structure do NOT matter. 3. The rules are short, concise, single sentences. Please number them from 1-num and do not say anything else
Online Inference with Hard Rules {spverbatim} Please list num possible rules about the att att_choices. Example 1: Structure: blue, blue Simple rules (Orders do NOT matter): 1. There is a blue block 2. There is no green block 3. All blocks are blue Task 1: x Simple rules (Orders do NOT matter):
Batch Inferece with Fuzzy Rules {spverbatim} Given the following structures described with att_summary of blocks in the structures: text_c Please list num possible rules about the attributes in a structure that differentiate the good structures from the bad structures. Keep in mind that 1. All bad structures must violate the rules. 2. Orders of blocks in a structure do NOT matter. 3. Do NOT propose ”negative rules” such as ”there is no green block”. 4. The rules are short, concise, single sentences. Please number them from 1-num and do not say anything else
Online Inference with Fuzzy Rules {spverbatim} Please list num possible rules about the att. Example 1: Structure: blue, blue Simple rules (Orders do NOT matter): 1. There is a blue block 2. All blocks are blue Do NOT propose ”negative rules” such as ”there is no green block”. Do NOT propose rules with quantifier such as ”there are two blue blocks” Task 1: x Simple rules (Orders do NOT matter):
Table 1: Prompts used for the proposer $q(h|x,y)$ of each method
{spverbatim} A structure has one or more blocks. Each block should contain the following attributes: color (blue/red/green) size (small/medium/large) orientation (upright/left/right/strange) groundedness (grounded/ungrounded), touching (which other blocks they do and do not touch). Example of rule modifications: Quantifier change: ’There must be a green block’ -¿ ’There are two green blocks’ Additional attribute: ’There must be a green block’ -¿ ’There must be a green block that is upright’ Attribute change: ’There must be a green block’ -¿ ’There must be a blue block’ These modifications are ”local”: only one attribute/quantifier is changed or added for each modification. Please modify the rule ’h’. Generate num rules for each type of modification (Quantifier change, Additional attribute, Attribute change) so that the following structure is text_y a good structure: x Note that the number of the blocks do not matter. Make the format a numbered list (1., 2., …, 15.) Remember that the new rules should be a ”local” modification from the rule ’h’. Do not use attribute values that are not mentioned earlier. Do not say anything other than the modified rules.
Table 2: Prompts used for the local proposer $q(h_{evolved}|h_{original},x,y)$ of online inferece methods
{spverbatim} A structure has one or more blocks. Each block should contain the following attributes: color (blue/red/green) size (small/medium/large) orientation (upright/left/right/strange) groundedness (grounded/ungrounded), touching (which other blocks they do and do not touch). Consider the following rule: ’h’ Given a structure, the output is yes if it follows the rule (or is a good structure) and no if it does not (or is a bad structure) The given rule gives incorrect output for the following structures: feedback Based on the given rule, generate num new refined rules that fix the outputs for all mentioned structures. The new rules may involve any of the mentioned attributes (color, size, orientation, grounded, touching). Please number them from 1-num and do not say anything else
Table 3: Prompts used for refinement (in batch inference with refinement with hard rules)
Initial prompt {spverbatim} You are playing an inductive game with me. I’ll be the moderator, and your task is to figure out the secret rule that I know by coming up with a structure of blocks to ask me whether it conforms with the secret rule or not. The structure has one of more blocks. Each block should contain the following attributes: color (blue/red/green) size (small/medium/large) orientation (upright/left/right/strange) groundedness (grounded/ungrounded), touching (which other blocks they do and do not touch). To give you a start, I’ll describe one structure that follows the rule: text_c Give a very short summary on what you currently think the secret rule is.
Follow-up prompt The verdict on whether the queried structure follows the rule is {verdict}. Give a very short summary on what you currently think the secret rule is.
Active learning (experiment) prompt Give one structure you want to test whether it follows the secret rule or not. Do not include anything other than the structure.
Table 4: Prompts used for vanilla, direct LLM method